# Real Noncommutative Convexity II: Extremality and nc convex functions

### v1

@@ 4. Noncommutative convex functions
${\rm Epi}_{m}(f)=\{(x,\alpha)\in K_{m}\otimes M_{n}(M_{m}):x\in K_{m},\alpha\geq f(x)\}.$
[…]
@@
This is simply a remark in [14] without proof. We add a proof, partly to serve as a model for the later more complicated multivalued nc function case. For $f\in M_{n}(B(K))$, we have that $f$ is graded, and so $\text{Epi}(f)$ is graded. Also $f$ preserves direct sums, which results in $\text{Epi}(f)$ being closed under direct sums. Indeed, let $(x_{i},\beta_{i})\in\text{Epi}_{m_{i}}(f)$, and let $\alpha_{i}\in M_{k,m_{i}}(\mathbb{R})$ be a family of isometries such that $\sum_{i}\alpha_{i}\alpha_{i}^{\mathsf{T}}=I_{k}$. We need to check that
@@
and because $f$ is closed under direct sums we get
@@
Remark. Remark 7.2.27.2.27.2.2 in [14] holds in the real case. Also, the Hansen-Pedersen type application in [14, Proposition 7.2.3] holds in the real case, showing that $f$ is real nc convex if and only if it is classically convex at every ‘level’.
@@
Let $f\in M_{n}(B(K))$, then $f$ is real nc convex (resp. lowersemicontinuous) if and only if $f_{c}=\Psi(f+i0)$ is complex convex (resp. lowersemicontinuous) as an element of $M_{n}(B(K_{c}))$
@@
Indeed $\mu_{c},\nu_{c}:C(K_{c})\to M_{m}(\mathbb{C})$ are ucp. Claim: $\mu_{c}(f+ig)=\nu_{c}(f+ig)$ for all $n<\infty$ and $f+ig\in M_{n}(C(K))=M_{n}(C(K_{c}))$ nc convex. If the Claim is true then by the complex case (Proposition 7.2.8 in [14]), $\mu_{c}=\nu_{c}$ so that $\mu=\nu$ as desired. Applying the map $c$, we see that $\mu_{c}(f+ig)=\nu_{c}(f+ig)$ if $\mu_{2}(c(f,g))=\nu_{2}(c(f,g))$, and the latter holds by hypothesis since $c(f,g)$ is nc convex by the lemma. ∎
@@ 5. Convex multivalued functions and convex envelopes
We show that if $F$ is bounded then $F_{c}$ is too. Indeed let $\lambda>0$ be such that for every $\beta\in F(x)$ there is an $\alpha\in F(x)$ with $\alpha\leq\beta$ and $\|\alpha\leq\lambda$. If $\beta\in F_{c}(x+iy)$ then there exists $\gamma\in F(c(x+iy))$ with $\beta=u^{*}\gamma u$. Choose $\alpha\in F(c(x+iy))$ with $\alpha\leq\gamma$ and $\|\alpha\|\leq\lambda$. Then $u^{*}\alpha u\in F(c(x+iy))$ with $u^{*}\alpha u\leq\beta$ and $\|u^{*}\alpha u\|\leq\|\alpha\|\leq\lambda$.
@@
[…]
@@
[…]
@@
Fix $x\in K_{m}$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$. If $z+\alpha\in\pi(G)(x)+M_{n}(M_{m}(\mathbb{R}))^{+}$ for $z=\text{Re }(y)$ where $y\in G(x)$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$, then $y+\alpha\in G(x)$ and so $\text{Re }(y)+\alpha\in\pi(G)(x)$. Thus $\pi(G)$ is upward directed. To show boundedness, let $x\in K$ and $\beta\in\pi(G)(x)$ so that $\beta=\text{Re }y$ for some $y\in G(x+i0)$. Because $G$ is bounded there exists a $\lambda>0$ and $\alpha\in G(x+i0)$ with $\alpha\leq y$ and $||\alpha||\leq\lambda$. Taking the Re of the inequality and using that Re is a contraction gives $\text{Re }\alpha\leq\text{Re }y=\beta$ and $||\text{Re }\alpha||\leq\lambda$. Thus, $\lambda$ is a bound for the function $\pi(G)$.
@@
They point out that this is not an nc multivalued function in the technical sense.
@@ 6. Appendix by Travis Russell on nc extreme points
If $\phi:\mathcal{S}\to B(H)$ and $\psi:\mathcal{S}\to B(K)$ are ucp maps on a real operator system, then $\phi\leq\psi$ below means that $H\subseteq K$ and $\phi$ is the restriction of $\psi$ to $H$ (i.e. $\psi$ is a dilation of the map $\phi$). As usual $\phi$ is maximal if every such $\psi$ is a trivial dilation as defined in Section 2. Let $\varphi:\mathcal{S}\to B(H)$ be a ucp map, and let $F\subset S\times H$. We say that $\varphi$ is maximal on $F$ if whenever $\psi\geq\varphi$, then $\psi(x)h=\varphi(x)h$ for all $(x,h)\in F$. If $\varphi:\mathcal{S}\to B(H)$ is maximal on $S\times H$, then it is maximal as defined above.
@@
More detailed proof of Theorem 2.20. As said below Theorem 2.20, with a few more specifics, we will use below that various properties of irreducible representations and corresponding GNS representations from [17], hold in the real case. We amplify Davidson and Kennedy’s argument, in part to see that there are no hidden real vs. complex issues. We will also use the following modified version of [14, Proposition 2.2.9]: Let $K$ be a closed nc-convex set over a dual operator space $E$. Suppose there is a net of contractions $\{\alpha_{i}\in M_{n,n_{i}}\}$ and a net $\{x_{i}\in K_{n_{i}}\}$ such that $\lim\alpha_{i}\alpha_{i}^{*}=I_{n}$ and such that $\lim\alpha_{i}x_{i}\alpha_{i}^{*}=x\in M_{n}(E)$. Then $x\in K_{n}$. To see this, let $\alpha_{i}=U_{i}P_{i}$ be a polar decomposition for $\alpha_{i}$, where $U_{i}\in M_{n,n_{i}}$ is a partial isometry and $P_{i}\in M_{n_{i}}$ is a positive operator. Let $\beta_{i}=U_{i}(I_{n_{i}}-P_{i}^{2})^{1/2}\in M_{n,n_{i}}$. Then $\alpha_{i}\alpha_{i}^{*}+\beta_{i}\beta_{i}^{*}=U_{i}U_{i}^{*}$, which is a projection in $M_{n}$. Let $m_{i}\leq n$ be a cardinal and $V_{i}\in M_{n,m_{i}}$ be a partial isometry such that $V_{i}V_{i}^{*}+U_{i}U_{i}^{*}=I_{n}$. Then because $\alpha_{i}\alpha_{i}^{*}\to I_{n}$ in $M_{n}(E)$, $\beta_{i}\beta_{i}^{*}\to 0$ and $V_{i}V_{i}^{*}\to 0$ in $M_{n}(E)$. Fix $y\in K_{1}$. Since $K$ is closed under nc-convex sets, $\alpha_{i}x_{i}\alpha_{i}^{*}+\beta_{i}(y\otimes I_{n_{i}})\beta_{i}^{*}+V_{i}(y\otimes I_{m_{i}})V_{i}^{*}\in K_{n}$ for every $i$. Since
@@
Since $\delta_{y}\in Z$ and since $Z$ is the closed nc convex hull of $\{\delta_{x}:x\in X\}$, if $a\in\cap_{x\in X}\ker(\delta_{x})$, then $a\in\ker(\delta_{y})$. So $\cap_{x\in X}\ker(\delta_{x})\subseteq\ker(\delta_{y})$. Consequently, $\cap_{x\in X}\ker(\delta_{x}\otimes id_{n})\subseteq\ker(\delta_{y}\otimes id_{n})$. Because the state $\psi$ factors through $\delta_{y}\otimes id_{n}$ and hence vanishes on $\ker(\delta_{y}\otimes id_{n})$, the real case of Dixmier’s Prop 3.4.2(ii) implies that $\psi$ is a weak-* limit of pure states $\psi_{x_{i}}:M_{n}(C(K))\to\mathbb{C}$ factoring through representations $\delta_{x_{i}}\otimes id_{n}$, i.e. $\psi_{x_{i}}(z)=\langle(\delta_{x_{i}}\otimes id)(z)h_{i},h_{i}\rangle$ for some unit vector $h_{i}\in H_{i}^{n}$, where $\delta_{x_{i}}:C(K)\to B(H_{i})$.

### v2

@@ 4. Noncommutative convex functions
${\rm Epi}_{m}(f)=\{(x,\alpha)\in K_{m}\times M_{n}(M_{m}):x\in K_{m},\alpha\geq f(x)\}.$
Davidson and Kennedy define $f$ to be (nc) convex if Epi(f)(f) is an nc convex set, and to be (nc) lowersemicontinuous (or ‘lsc’) if $\text{Epi}(f)$ is closed.
@@
In the complex case this is simply a remark in [14] without proof. We add a proof, partly to serve as a model for the later more complicated multivalued nc function case. For $f\in M_{n}(B(K))$, we have that $f$ is graded, and so $\text{Epi}(f)$ is graded. Also $f$ preserves direct sums, which results in $\text{Epi}(f)$ being closed under direct sums. Indeed, let $(x_{i},\beta_{i})\in\text{Epi}_{m_{i}}(f)$, and let $\alpha_{i}\in M_{k,m_{i}}(\mathbb{R})$ be a family of isometries such that $\sum_{i}\alpha_{i}\alpha_{i}^{\mathsf{T}}=I_{k}$. We need to check that
@@
and because $f$ respects direct sums we get
@@
Remark. Remark $7.2.2$ in [14] holds in the real case with the same proof. Similarly, the Hansen-Pedersen type application in [14, Proposition 7.2.3] holds in the real case, showing that $f$ is real nc convex if and only if it is classically convex at every ‘level’.
@@
Let $f\in M_{n}(B(K))$, then $f$ is real nc convex (resp. lowersemicontinuous) if and only if $f_{c}=\Psi(f+i0)$ is complex convex (resp. lowersemicontinuous) as an element of $M_{n}(B(K_{c}))$.
@@
Indeed $\mu_{c},\nu_{c}:C(K_{c})\to M_{m}(\mathbb{C})$ are ucp. Claim: $\mu_{c}(f+ig)=\nu_{c}(f+ig)$ for all $n<\infty$ and $f+ig\in M_{n}(C(K))=M_{n}(C(K_{c}))$ nc convex. If the Claim is true then by the complex case (Proposition 7.2.8 in [14]), $\mu_{c}=\nu_{c}$ so that $\mu=\nu$ as desired. Applying the map $c$, we see that $\mu_{c}(f+ig)=\nu_{c}(f+ig)$ if $\mu_{2}(c(f,g))=\nu_{2}(c(f,g))$, and the latter holds by hypothesis since $c(f,g)$ is nc convex by the previous lemma. ∎
@@ 5. Convex multivalued functions and convex envelopes
We show that if $F$ is bounded then $F_{c}$ is too. Indeed let $\lambda>0$ be such that for every $\beta\in F(x)$ there is an $\alpha\in F(x)$ with $\alpha\leq\beta$ and $\|\alpha\|\leq\lambda$. If $\beta\in F_{c}(x+iy)$ then there exists $\gamma\in F(c(x+iy))$ with $\beta=u^{*}\gamma u$. Choose $\alpha\in F(c(x+iy))$ with $\alpha\leq\gamma$ and $\|\alpha\|\leq\lambda$. Then $u^{*}\alpha u\in F(c(x+iy))$ with $u^{*}\alpha u\leq\beta$ and $\|u^{*}\alpha u\|\leq\|\alpha\|\leq\lambda$.
@@
for all $x\in K_{m}$ and isometries $\alpha\in M_{k,m}(\mathbb{R})$. Indeed clearly Graph(F)(F) is graded. Suppose that $(x,\beta)\in{\rm Graph}_{m}(F)$ and $\alpha\in M_{m,k}$ is an isometry. We have $\beta\in F(x)$ and $\alpha^{\mathsf{T}}x\alpha\in K_{k}$. Using (2) gives
@@
and so $\alpha^{\mathsf{T}}(x,\beta)\alpha\in{\rm Graph}(F)$. Conversely, if ${\rm Graph}(F)$ is a nc convex set then for all $x\in K_{m}$ and isometries $\alpha\in M_{k,m}$ we have $(x,F(x))\in{\rm Graph}_{m}(F)$ and therefore $\big{(}\alpha^{\mathsf{T}}x\alpha,(1_{n}\otimes\alpha^{\mathsf{T}})F(x)(1\otimes\alpha)\big{)}\in{\rm Graph}(F)$ which gives (2). A similar argument to that in Lemma 4.1 shows that Graph(F)(F) is closed under direct sums. So $F$ is nc convex.
@@
Fix $x\in K_{m}$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$. If $z+\alpha\in\pi(G)(x)+M_{n}(M_{m}(\mathbb{R}))^{+}$ for $z=\text{Re }(y)$ where $y\in G(x)$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$, then $y+\alpha\in G(x)$ and so $\text{Re }(y)+\alpha\in\pi(G)(x)$. Thus $\pi(G)$ is upward directed. To show boundedness, let $x\in K$ and $\beta\in\pi(G)(x)$ so that $\beta=\text{Re }y$ for some $y\in G(x+i0)$. Because $G$ is bounded there exists a $\lambda>0$ and $\alpha\in G(x+i0)$ with $\alpha\leq y$ and $\alpha||\leq\lambda$. Taking the Re of the inequality and using that Re is a contraction gives $\text{Re }\alpha\leq\text{Re }y=\beta$ and $\text{Re }\alpha||\leq\lambda$. Thus, $\lambda$ is a bound for the function $\pi(G)$.
@@
They point out that this is not a nc multivalued function in the technical sense.
@@ 6. Appendix by Travis Russell on nc extreme points
If $\varphi:\mathcal{S}\to B(H)$ and $\psi:\mathcal{S}\to B(K)$ are ucp maps on a real operator system, then $\varphi\leq\psi$ below means that $H\subseteq K$ and $\varphi$ is the restriction of $\psi$ to $H$ (i.e. $\psi$ is a dilation of the map $\varphi$). As usual $\varphi$ is maximal if every such $\psi$ is a trivial dilation as defined in Section 2. Let $\varphi:\mathcal{S}\to B(H)$ be a ucp map, and let $F\subset S\times H$. We say that $\varphi$ is maximal on $F$ if whenever $\psi\geq\varphi$, then $\psi(x)h=\varphi(x)h$ for all $(x,h)\in F$. If $\varphi:\mathcal{S}\to B(H)$ is maximal on $S\times H$, then it is maximal as defined above.
@@
More detailed proof of Theorem 2.20. As said (together with a few more specifics) below Theorem 2.20, we will use below that various properties of irreducible representations and corresponding GNS representations from [17], hold in the real case. We amplify Davidson and Kennedy’s argument, in part to see that there are no hidden real vs. complex issues. We will also use the following modified version of [14, Proposition 2.2.9]: Let $K$ be a closed nc-convex set over a dual operator space $E$. Suppose there is a net of contractions $\{\alpha_{i}\in M_{n,n_{i}}\}$ and a net $\{x_{i}\in K_{n_{i}}\}$ such that $\lim\alpha_{i}\alpha_{i}^{*}=I_{n}$ and such that $\lim\alpha_{i}x_{i}\alpha_{i}^{*}=x\in M_{n}(E)$. Then $x\in K_{n}$. To see this, let $\alpha_{i}=U_{i}P_{i}$ be a polar decomposition for $\alpha_{i}$, where $U_{i}\in M_{n,n_{i}}$ is a partial isometry and $P_{i}\in M_{n_{i}}$ is a positive operator. Let $\beta_{i}=U_{i}(I_{n_{i}}-P_{i}^{2})^{1/2}\in M_{n,n_{i}}$. Then $\alpha_{i}\alpha_{i}^{*}+\beta_{i}\beta_{i}^{*}=U_{i}U_{i}^{*}$, which is a projection in $M_{n}$. Let $m_{i}\leq n$ be a cardinal and $V_{i}\in M_{n,m_{i}}$ be a partial isometry such that $V_{i}V_{i}^{*}+U_{i}U_{i}^{*}=I_{n}$. Then because $\alpha_{i}\alpha_{i}^{*}\to I_{n}$ in $M_{n}(E)$, $\beta_{i}\beta_{i}^{*}\to 0$ and $V_{i}V_{i}^{*}\to 0$ in $M_{n}(E)$. Fix $y\in K_{1}$. Since $K$ is closed under nc-convex sets, $\alpha_{i}x_{i}\alpha_{i}^{*}+\beta_{i}(y\otimes I_{n_{i}})\beta_{i}^{*}+V_{i}(y\otimes I_{m_{i}})V_{i}^{*}\in K_{n}$ for every $i$. Since
@@
Since $\delta_{y}\in Z$ and since $Z$ is the closed nc convex hull of $\{\delta_{x}:x\in X\}$, if $a\in\cap_{x\in X}\ker(\delta_{x})$, then $a\in\ker(\delta_{y})$. So $\cap_{x\in X}\ker(\delta_{x})\subseteq\ker(\delta_{y})$. Consequently, $\cap_{x\in X}\ker(\delta_{x}\otimes id_{n})\subseteq\ker(\delta_{y}\otimes id_{n})$. Because the state $\psi$ factors through $\delta_{y}\otimes id_{n}$ and hence vanishes on $\ker(\delta_{y}\otimes id_{n})$, the real case of Dixmier’s Proposition 3.4.2(ii) implies that $\psi$ is a weak-* limit of pure states $\psi_{x_{i}}:M_{n}(C(K))\to\mathbb{C}$ factoring through representations $\delta_{x_{i}}\otimes id_{n}$, i.e. $\psi_{x_{i}}(z)=\langle(\delta_{x_{i}}\otimes id)(z)h_{i},h_{i}\rangle$ for some unit vector $h_{i}\in H_{i}^{n}$, where $\delta_{x_{i}}:C(K)\to B(H_{i})$.